# Spectra and transport of probability and energy densities in a $\mathcal{PT}$-symmetric square well with a delta-function potential

### v1

@@ I Introduction
Since its introduction by Bender and Boettcher Bender and Boettcher (1998), $\mathcal{P}\mathcal{T}$-symmetric quantum mechanics has attracted much interest due to the phenomena and applications that it can describe, see, for example Bagarello et al. (2016); El-Ganainy et al. (2018); Christodoulides et al. (2018); Bender (2019) and references therein. One of the most studied aspects of non-Hermitian $\mathcal{PT}$-symmetric Hamiltonians is the appearance of exceptional points as a parameter of the system is varied. At such points, the $\mathcal{PT}$ symmetry of the solutions breaks, the eigenvalues become complex Zyablovsky et al. (2014), and two or more eigenvectors coalescence. As a result, the spectral decomposition of operators fails, and the Hamiltonian becomes defective Kato (2013); Heiss (2012); Bender (2019).
@@
To gain insight on the phenomenology that can arise in this kind of systems, it is desirable to have analytical models in which a system with a $\mathcal{P}\mathcal{T}$-symmetric potential can be solved analytically (see for instance Moiseyev (2011) and references therein for a general reference on non-Hermitian systems). For instance Lévai (2008) addresses central potentials while references below address one-dimensional square well and step potentials. However, only a few $\mathcal{P}\mathcal{T}$-symmetric potentials are solvable Bender (2019), perhaps the most explored have been one-dimensional square wells in their distinct versions. The first system of this type was a complex extension of the finite potential well where its bound states were calculated Znojil (2001). Due to the simplicity of this kind of systems, they have been studied in frameworks such as pseudo-Hermitian quantum mechanics Mostafazadeh and Batal (2004) as well as from the perspective of supersymmetry BAGCHI et al. (2006), while other studies have focused on aspects such as the breaking of the $\mathcal{P}\mathcal{T}$-symmetry in a step potential ZNOJIL and LÉVAI (2001). Later, modifications with periodic boundary conditions Jakubský and Znojil (2004), discrete systems Znojil (2006a), two coupled $\mathcal{P}\mathcal{T}$-symmetric square wells Znojil (2006b) and spatially antisymmetric Quesne et al. (2005) versions also have been investigated. Surprisingly, $\mathcal{P}\mathcal{T}-$symmetric finite wells in an infinite interval, whose Hermitian form is traditionally addressed in almost any standard mechanics courseSprung et al. (1996), have barely been studied Lévai and Kovács (2018). This model is interesting because it presents scattering states as well as an infinite set of bound state solutions Lévai and Kovács (2018). In this work we consider a generalization of this model that includes a delta-potential at the origin, which gives rise to the appearance of exceptional points as well.
@@
Regarding the scattering aspects of the non-Hermitian $\mathcal{P}\mathcal{T}$-symmetric potential, it turns out that the transmission and reflection coefficients do not obey the usual unitary rule. Instead, other “conservation rules” have been established Ge et al. (2012); Lin et al. (2011); Cannata et al. (2007); Ahmed (2001); Mostafazadeh (2014). Also, it has been reported that these coefficients present anomalous behaviour, i.e. they can take values greater than one Ahmed (2004), or even be infinite at certain values in the continuous spectrum Mostafazadeh (2009); Chaos-Cador and García-Calderón (2013).
In this work, we study an extension of the $(\mathcal{P}\mathcal{T}$)-symmetric square potential well presented by Znojil Znojil (2005) and Lévai Lévai and Kovács (2018). Yet, unlike those systems, ours exhibits a broken $\mathcal{P}\mathcal{T}$-symmetry phase. Our system consists of a finite $\mathcal{PT}$-symmetric complex potential well and a delta-potential at the origin modulated by a (real) coupling constant $\lambda$. This enables the system to exhibit a complexification of its eigenvalues as a function of this parameter, thereby displaying the characteristics specific to this phenomenon, such as the appearance of exceptional points. In addition to its spectral properties, we take also analyze its transport properties. For this, we compute the associated density currents for bound and scattering states, as well as the associated energy densities.
@@ II.1 Discrete spectrum and bound states
The exceptional points appear at different values of $\Lambda$ as $k$ grows. We can approximate their behavior when $k\gg v_{0}$ as follows. Expanding $\alpha_{R}$ and $\alpha_{I}$ for large values of $k$ we have
$\alpha_{R}\sim\frac{v_{I}}{2k},$ (13)
$\alpha_{I}\sim k\left(1-\frac{v_{0}}{2k^{2}}\right).$ (14)
Substituting in Equation (12) and expanding to leading order in $k$ we have
@@
This equation has solution if we take $\Lambda\sim\chi/k$, where $\chi$ is constant
@@
$\chi<|\sqrt{v_{0}^{2}+v_{I}^{2}}|.$ (18)
Hence the curve that describes the location of the exceptional points for $k\gg 0$ is approximately $\kappa(\Lambda)=\sqrt{v_{0}^{2}+v_{I}^{2}}/\Lambda$. We show this curve in figure (1), dashed purple line.
Figure 1: The blue curves represent the pseudo-momentum $k$ for the bound states as a function of $\lambda$. These curves end at exceptional points at values approximated by $\kappa(\Lambda)$ in Eq.(18) (purple line), for states with energies greater than $\sqrt{v_{0}}$ (red dashed line). Additionally, we can see the solutions for $\Lambda=0$ (gray lines). In the insets one can observe the complexification of the spectrum of the first six states as a function of the parameter $\Lambda$, here $k_{R}$ and $k_{I}$ denote the real and imaginary parts of $k$.
@@
Figure 2: Density of the bound states for cases when $k^{2}<V_{0}$ a) and $k^{2}>V_{0}$ b) for some values of $v_{I}$. As the magnitude of the imaginary potential decreases, the bound states with $k^{2}>v_{0}$ become more extended and their magnitude decreases accordingly.
@@ II.2 Scattering States
In what follows we describe the properties of scattering states that occur in this system Cannata et al. (2007); Mostafazadeh (2018) . The specific scattering processes can be seen schematically in Fig. 3.
Figure 3: Scattering process. a) Left-to-right incidence and b) right-to-left incidence. The figure also illustrates the meaning of the coefficients $r_{+}$, $r_{-}$, $t_{+}$ and $t_{-}$.
The solution of the Schrödinger equation for the left-to-right scattering case can be obtained making the coefficient $B_{1}=0$ in Eq. (7), thus
@@
The scattering solutions are not $\mathcal{P}\mathcal{T}-$symmetric, and due to the presence of a source and a sink, the relation $|\mathbf{r}_{+}|^{2}+|\mathbf{t}_{-}|^{2}=1$ is not satisfied. Instead we must consider the inverse process, i.e. the right-to-left scattering with wave function given by (c.f. Eq. (7))
@@
Figure 4: a) Pseudo-transmission $T_{+/-}$, b) Left pseudo-reflection $R_{+}$, c) Right pseudo-reflection $R_{-}$ d) Module of the product of left and right pseudo-reflections as a function of $k$ for the parameter $v_{I}=10$, $v_{I}=10$ and $b=1$.
From these equations we can compute the pseudo-transmissions $T_{\pm}=|\mathbf{t_{\pm}}|^{2}$ and pseudo-reflections $R_{\pm}=|\mathbf{t_{\pm}}|^{2}$, which are shown in Fig. 4. We note that the the left and right transmission are equal, $|\mathbf{t}_{+}(k)|^{2}=|\mathbf{t}_{-}(k)|^{2}$. Actually, the pseudo-transmissions and pseudo-reflection can show anomalous behavior, taking values greater than one as can be seen in Fig. 4. On the other hand, the left reflectivity ($R_{+}$) vanishes at the bound states as expected. This can be observed in Fig. 4 where we recover the spectrum plot in the $(\Lambda,k)$-plane. The figure also shows that there is a set of discrete values of $k$ at which $R_{-}$ vanishes, which correspond to non integrable reflection-less states.
@@ III Transport properties
In this section we focus on the analysis of the transport properties of the system. We discuss briefly how different definitions for the energy density yield different continuity equations, then we analyze both definitions in the context of our specific problem.
@@ III.1 Probability density, energy density and their continuity equations
The particle is described by the wavefunction $\psi(x,t)$ for which the probability density is $\rho_{d}(x,t)=|\psi(x,t)|^{2}$. This probability density satisfies the (standard) continuity equation
@@
accounts for the source/drain terms. Under this perspective, the imaginary part of the potential is interpreted as giving rise to a source or sink of probability Molinàs-Mata and Molinàs-Mata (1996). However, it is important to mention that a self-consistent conservation law has been proposed for the norm defined by the $\mathcal{P}\mathcal{T}$ inner product BAGCHI et al. (2001); Japaridze (2002). However, this norm usually presents certain drawbacks, so for the purposes of this work, we will not adopt it. Instead, we will approach the transport properties from the perspective of the continuity equation (36).
Another quantity that satisfies a continuity equation in quantum mechanics, and is less explored, is the energy density Mathews (1974); Ludovico et al. (2014); Arvizu et al. (2023). Unlike the definition for the probability density, the definition for the energy density in quantum mechanics is not unique, different definitions yield different results and therefore different interpretations of the “flux” of energy in a quantum system. An initial attempt to define the energy density could be simply as
@@
so $V_{I}$ does not contribute to the expectation value of the energy. Instead, we note that the integrand of the previous equation is proportional to $Q_{d}(x,t)$ Molinàs-Mata and Molinàs-Mata (1996); Ahmed (2001); Wahlstrand et al. (2014), thus, equation (40) can be interpreted as saying that in $\mathcal{PT}-$symmetric systems gain and loss of density is globally balanced Weigert (2004); Ortega et al. (2020).
The condition (40) allows us to obtain a first acceptable definition of the energy density ($\rho_{1}^{E}$) for the $\mathcal{PT}-$symmetric phase as
@@
This definition yields real density, but has the drawback of not being necessarily positive even for non-negative potentials. Alternatively, another definition of energy density is
$\rho^{E}_{2}(x,t)=\frac{\hbar^{2}}{2m}\left|\frac{\partial}{\partial x}\psi(x,t)\right|^{2}+V_{R}(x)|\psi(x,t)|^{2},$ (42)
as can be checked by integrating by parts and using that the wave function vanishes at infinity.
@@
Now we analyze first Sec. III.1.1 the probability and energy transport properties of the bound states. Subsequently in Sec. III.1.2, we apply this analysis for the scattering states.
@@ III.1.1 Bound states
where $|C_{1}|^{2}$ is given in Eq.(22). We can see in this expression that the flux on the side of the source ($x\leq b$) increases as it approaches the edge of the well. Once there, it remains constant across the well, and it decreases on the other side of the well.
@@
It is interesting to see how the potential induces a discontinuity in the energy density $\rho_{2}^{E}$, while the behavior of $\rho_{1}^{E}$ is the same as that of the probability density.

### v2

@@ I Introduction
Since its introduction by Bender and Boettcher Bender and Boettcher (1998), $\mathcal{P}\mathcal{T}$-symmetric quantum mechanics has attracted much interest due to the phenomena and applications that it can describe, see, for example Bagarello et al. (2016); El-Ganainy et al. (2018); Christodoulides et al. (2018); Bender (2019) and references therein. One of the most studied aspects of non-Hermitian $\mathcal{PT}$-symmetric Hamiltonians is the appearance of exceptional points as a parameter of the system is varied. At such points, the $\mathcal{PT}$ symmetry of the solutions breaks, the eigenvalues become complex Zyablovsky et al. (2014), and two or more eigenvectors coalesce. As a result, the spectral decomposition of operators fails, and the Hamiltonian becomes defective Kato (2013); Heiss (2012); Bender (2019).
@@
To gain insight on the phenomenology that can arise in this kind of systems, it is desirable to have analytical models in which a system with a $\mathcal{P}\mathcal{T}$-symmetric potential can be solved analytically (see for example Moiseyev (2011) and references therein, for a general overview on non-Hermitian systems). For instance  Lévai (2008) addresses central potentials while references below address one-dimensional square well and step potentials. However, only a few $\mathcal{P}\mathcal{T}$-symmetric potentials are solvable Bender (2019), perhaps the most explored have been one-dimensional square wells in their distinct versions. The first system of this type was a complex extension of the finite potential well where its bound states were calculated Znojil (2001). Due to the simplicity of this kind of systems, they have been studied in frameworks such as pseudo-Hermitian quantum mechanics Mostafazadeh and Batal (2004) as well as from the perspective of supersymmetry  BAGCHI et al. (2006), while other studies have focused on aspects such as the breaking of the $\mathcal{P}\mathcal{T}$-symmetry in a step potential ZNOJIL and LÉVAI (2001). Later, modifications with periodic boundary conditions Jakubský and Znojil (2004), discrete systems Znojil (2006a); Barashenkov et al. (2013), two coupled $\mathcal{P}\mathcal{T}$-symmetric square wells  Znojil (2006b) and spatially antisymmetric Quesne et al. (2005) versions also have been investigated. Surprisingly, $\mathcal{P}\mathcal{T}-$symmetric finite wells in an infinite interval, whose Hermitian form is traditionally addressed in almost any standard mechanics course Sprung et al. (1996), have barely been studied Lévai and Kovács (2018). This model is interesting because it presents scattering states as well as an infinite set of bound state solutions Lévai and Kovács (2018). On the other hand, other solvable models related to $\mathcal{PT}-$symmetric square wells are systems containing delta potentials Znojil and Jakubský (2005); Uncu and Demiralp (2006); Barashenkov and Zezyulin (2016); Kovács and Lévai (2017); Mostafazadeh (2018). The mechanisms of the level crossings and of the $\mathcal{PT}$-symmetry breaking in these systems, as well as their scattering properties, have been widely studied Znojil and Jakubský (2005); Zhang and Gong (2017), and applied in the context of Bose-Einstein condensation Dast et al. (2013); Abt et al. (2015); Cartarius and Wunner (2012).
@@
Regarding the scattering aspects of the non-Hermitian $\mathcal{P}\mathcal{T}$-symmetric potential, it turns out that the transmission and reflection coefficients do not obey the usual unitary rule. Instead, other “conservation rules” have been established Ge et al. (2012); Lin et al. (2011); Cannata et al. (2007); Ahmed (2001a); Mostafazadeh (2014), which we discuss in detail in Section 3. Also, it has been reported that these coefficients may present anomalous behaviour, i.e. they can take values greater than one Ahmed (2004), or even be infinite at certain values in the continuous spectrum Mostafazadeh (2009); Chaos-Cador and García-Calderón (2013).
In this work, we study an extension of the $\mathcal{P}\mathcal{T}$-symmetric square potential well presented by Znojil Znojil (2005) and Lévai Lévai and Kovács (2018). Yet, unlike those systems, ours exhibits a broken $\mathcal{P}\mathcal{T}$-symmetry phase. Our system consists of a finite $\mathcal{PT}$-symmetric complex potential well and a delta-potential at the origin modulated by a (real) coupling constant $\lambda$. This enables the system to exhibit a complexification of its eigenvalues as a function of this parameter, thereby displaying the characteristics specific to this phenomenon, such as the appearance of exceptional points. Qualitatively similar phenomena occur in a related discrete Hamiltonian Barashenkov et al. (2013). Also, as mentioned above, a central feature of $\mathcal{PT}$-symmetric systems is the the presence of fluxes driven by the gain and loss of density and energy arising from the imaginary part of the potential, even in the $\mathcal{PT}$-symmetric stationary states. Thus, the characterization of the fluxes within these systems is crucial to understanding their behavior. In view of this, in addition to its spectral properties, we also analyze the transport properties of probability and energy within the system. Specifically, we compute and study the behavior of the density currents for bound and scattering states, as well as the associated energy densities and their respective fluxes. Also related to the transport properties of the system, we study the energy densities and energy fluxes. Since the system is solvable, we obtain closed form expressions for two energy densities definitions, and discuss their relevance in connection with the probability density and flux of probability.
@@ II.1 Discrete spectrum and bound states
The exceptional points appear at different values of $\Lambda$ as $k$ grows. We can approximate their behavior when $k\gg\sqrt{v_{0}}$ as follows (for a general reference on asymptotic expansions see Bender and Orszag (1999)). Expanding $\alpha_{R}$ and $\alpha_{I}$ from Eqs. (8) and (9) for large values of $k$, we find
$\alpha_{R}\approx\frac{v_{I}}{2k},$ (13)
$\alpha_{I}\approx k\left(1-\frac{v_{0}}{2k^{2}}\right).$ (14)
Substituting these expressions in Equation (12) and expanding to leading order in $k$ we have
@@
This equation has solution if we take $\Lambda\equiv\chi/k$, where $\chi$ is constant, thus
@@
$\chi\leq|\sqrt{v_{0}^{2}+v_{I}^{2}}|.$ (18)
Hence a curve that bounds the solutions of Eq.(11), including the exceptional points, for $k\gg\sqrt{v_{0}}$, is $\kappa(\Lambda)=\sqrt{v_{0}^{2}+v_{I}^{2}}/\Lambda$. We show this curve in Fig. 1 (dashed purple line).
Figure 1: The blue curves represent the pseudo-momentum $k$ for the bound states as a function of $\Lambda$ with $v_{0}=9$, $v_{I}=15$ and $b=1$. These curves end at exceptional points at values bounded by $\kappa(\Lambda)$ in Eq.(18) (purple dashed line) at large values of $k$ ($k>>\sqrt{v_{0}}$ solid green line). In the insets one can observe the complexification of the spectrum of the first eight states as a function of the parameter $\Lambda$, here $k_{R}$ and $k_{I}$ denote the real and imaginary parts of $k$.
@@
Figure 2: Density of the bound states for cases when $k^{2}<v_{0}$, a), and $k^{2}>v_{0}$, b), for some values of $v_{I}$ with $v_{0}=10$, $\Lambda=0.5$ and $b=1$. As the magnitude of the imaginary potential decreases, the bound states with $k^{2}>v_{0}$ become more extended and their magnitude decreases accordingly.
@@ II.2 Scattering States
In what follows we describe the properties of scattering states that occur in this system Cannata et al. (2007); Mostafazadeh (2018). The specific scattering processes can be seen schematically in Fig. 3. The complete description of these setups involves both the left-to-right and right-to-left processes, as the properties of the respective reflection and transmission coefficients are intrinsically related to the scattering direction Garmon et al. (2015); Mostafazadeh (2014), as we will show below.
Figure 3: Scattering process. a) Left-to-right incidence and b) right-to-left incidence. The figure also illustrates the reflection $r_{\pm}$ and transmission coefficients $t_{\pm}$ involved in the two scattering processes.
First we find the solution of the Schrödinger equation for the left-to-right scattering case can be obtained making the coefficient $B_{1}=0$ in Eq. (7), thus
@@
On the other hand, the right-to-left scattering is described by a wave function of the form (c.f. Eq. (7))
@@
Figure 4: a) Pseudo-transmission $T_{+/-}$, b) Left pseudo-reflection $R_{+}$, c) Right pseudo-reflection $R_{-}$ d) Module of the product of left and right pseudo-reflections as a function of $k$ for the parameter $v_{I}=10$, $v_{I}=10$ and $b=1$. The black vertical lines in b) indicate the solutions of Eq. (11). We can see how the zeros of Eq. (11) indicate where the reflection coefficient $R_{+}(k)$ vanishes, which, due to the logarithmic scale of the figure, can be seen as the curves tending to negative infinity.
From these results we can compute the pseudo-transmissions $T_{\pm}=|\mathbf{t_{\pm}}(k)|^{2}$ and pseudo-reflections $R_{\pm}=|\mathbf{r_{\pm}}(k)|^{2}$, which are shown in Fig. 4. We note that the left and right transmission are equal, $|\mathbf{t}_{+}(k)|^{2}=|\mathbf{t}_{-}(k)|^{2}$. Actually, the pseudo-transmissions and pseudo-reflection can show anomalous behavior, taking values greater than one as can be seen in Fig. 4. On the other hand, the left reflectivity ($R_{+}$) vanishes at the bound states, as expected. This can be observed in Fig. 4 or explicitly in the numerator of the expression for $\mathbf{r}_{+}(k)$ in Eq. (31). Indeed, when the left reflectivity coefficient is zero, the wave function Eq. (25) takes the form of the bound state wave functions given in Eq. (10). Fig. 4 also shows that there is a set of discrete values of $k$ at which $R_{-}$ vanishes, which correspond to non integrable reflection-less states, i.e. states for which the non vanishing terms in Eq. (28) diverge as $|x|\rightarrow\infty$.
@@ III Transport properties
In this section we focus on the analysis of the transport properties of the system. Specifically, we center our attention on the probability and energy densities and their respective fluxes. However, to do so, it is convenient to discuss briefly two different definitions for the energy density, as these yield different continuity equations, and only then analyze their transport properties in the context of our specific problem.
@@ III.1 Probability density, energy density and their continuity equations
The particle is described by the wavefunction $\psi(x,t)$ for which the probability density is $\rho_{d}(x,t)=|\psi(x,t)|^{2}$. This probability density satisfies the (standard) continuity equation, which can be derived by computing the time derivative of the probability density $\rho$ and the Schödinger equation, along with its complex conjugate, and taking into account the complex potential in Eq. (35). Several examples of such equation include for instance Molinàs-Mata and Molinàs-Mata (1996); Kočinac and Milanović (2008); Ernzerhof and Zhuang (2003); Elenewski and Chen (2014); Ahmed (2001b); Zhang et al. (2021). The resulting continuity equation is
@@
accounts for the source/drain terms. Thus, under this perspective, the imaginary part of the potential is interpreted as giving rise to a source or sink of probability Molinàs-Mata and Molinàs-Mata (1996). It is important to mention that a self-consistent conservation law has been proposed for the norm defined by the $\mathcal{P}\mathcal{T}$ inner product BAGCHI et al. (2001); Japaridze (2002). However, this norm usually presents certain drawbacks, so for the purposes of this work, we will not adopt it. Instead, we will approach the transport properties of the probability density from the perspective of the continuity equation (36).
Another quantity that satisfies a continuity equation in quantum mechanics, and is less explored, is the energy density Mathews (1974); Ludovico et al. (2014); Arvizu et al. (2023). Unlike the definition for the probability density, the definition for the energy density in quantum mechanics is not unique, different definitions yield different results and therefore may give different values for the flux of energy in a quantum system. An initial attempt to define the energy density could be simply as
@@
so $V_{I}$ does not contribute to the expectation value of the energy. Instead, we note that the integrand of the previous equation is proportional to $Q_{d}(x,t)$ Molinàs-Mata and Molinàs-Mata (1996); Ahmed (2001a); Wahlstrand et al. (2014), therefore, equation (40) can be interpreted as saying that in $\mathcal{PT}-$symmetric systems gain and loss of density is globally balanced Weigert (2004); Ortega et al. (2020). Thus, $\mathcal{PT}-$symmetric stationary states must have sustained currents carrying the probability density and energy density that are being input to where they are being removed.
Equation (40) allows us to obtain a first acceptable definition of the energy density ($\rho_{1}^{E}$) for the $\mathcal{PT}-$symmetric phase as
@@
This definition yields a real energy density, but has the drawback of not being necessarily positive even for non-negative potentials. Alternatively, another definition of energy density is
$\rho^{E}_{2}(x,t)=\frac{\hbar^{2}}{2m}\left|\frac{\partial}{\partial x}\psi(x,t)\right|^{2}+V_{R}(x)|\psi(x,t)|^{2}.$ (42)
This expression also integrates to the total expected energy of the system, as can be checked by integrating by parts and using that the wave function vanishes at infinity, and has a positive kinetic energy density. The physical relevance, of either expression is an open question Stepanyan and Allahverdyan (2024); Mathews (1974); Arvizu et al. (2023), so in what follows, we study the properties for both expressions.
@@
In Sec. III.1.1 we first analyze the probability and energy transport properties of the bound states. Subsequently, in Sec. III.1.2, we analyze the transport properties of the scattering states.
@@ III.1.1 Bound states
where $|C_{1}|^{2}$ is given in Eq.(22). We can see in this expression that the flux on the side of the source ($x\leq b$) increases as it approaches the edge of the well. Once there, it remains constant across the well, and it decreases on the other side of the well. Because of this, the flux inside the well can be viewed as the flux across the loss-gain interface, which depends on the parameters $v_{I}$ and $\Lambda$. These two parameters, together with $v_{0}$ and $b$, through Eq. (11), determine the values of $k$ corresponding to the spectrum of bound states and, consequently, the flux corresponding to each of these states. In the left panel of Fig. 5, the flux behaves in a manner analogous to $k$ (Fig. 1), when $v_{I}$, $v_{0}$ and $b$ are kept constant, while $\Lambda$ is varied. In this case, both $k$ and $J_{d}(0)$ corresponding to successive states either increase or decrease monotonously with $\Lambda$. These two values of $k$ and $J_{d}(0)$ eventually coalesce at the exceptional point  Barashenkov et al. (2016). In the right panel, the flux’s dependence on the parameter $v_{I}$ is shown, with $\Lambda$ and $v_{0}$ held constant. We observe that the flux is an increasing monotonic function of $v_{I}$. This is consistent with systems similar to ours, such as the $\mathcal{PT-}$symmetric square well without a delta function potential, that never shows a broken $\mathcal{PT-}$symmetry phase as a function $v_{I}$ Lévai and Kovács (2018), and there is no saturation of probability. In our system the symmetry breaking is due to the combination of $v_{I}$ and $\Lambda$, so the role attributable to the source in the systems studied in Barashenkov et al. (2016) can be attributed in principle to the role played by the delta potential as a barrier.
Figure 5: Plot of the flux inside the well for each of the bound states corresponding to each value of $\Lambda$ (with $v_{0}=9$ and $b=1$) . In a) in which we keep $v_{I}=15$ constant, we observe that the flux is alternating between increasing and decreasing with $\Lambda$, up to the exceptional point at which the states coalesce. On the other hand, in b) where we keep $\Lambda=0.5$ fixed, we see that $J_{d}(0)$ is always a monotonously increasing function of $v_{I}$.
@@
It is interesting to see that for the bound states, the energy density flux is the same for both definitions, while the energy densities are markedly different. The behavior of $\rho_{1}^{E}$ is essentially the same as that of the probability density, while the potential induces a discontinuity in the energy density $\rho_{2}^{E}$.